Ninety liters of pure maple syrup that costs $8.00 per liter are mixed with imitation maple syrup that costs $2.80 per liter. How much imitation maple syrup is needed to make a mixture to sell for $5.14 per liter?
Answer by checkley77(12844)

8*90+2.80x=5.14(90+x)
720+2.80x=462.6+5.14x
2.80x-5.14x=462.6-720
-2.34x=-257.4
x=-257.4/2.34
x=110 liters of 2.80 syrup will be required.
proof:
8*90+2.80*110=5.14(90+110)
720+308=5.14*200
1028=1028